# PRISM: A Robust Framework for Skill-based Meta-Reinforcement Learning with Noisy Demonstrations

### v2

@@ Abstract
Meta-reinforcement learning (Meta-RL) facilitates rapid adaptation to unseen tasks but faces challenges in long-horizon environments. Skill-based approaches tackle this by decomposing state-action sequences into reusable skills and employing hierarchical decision-making. However, these methods are highly susceptible to noisy offline demonstrations, resulting in unstable skill learning and degraded performance. To overcome this, we propose Prioritized Refinement for Skill-Based Meta-RL (PRISM), a robust framework that integrates exploration near noisy data to generate online trajectories and combines them with offline data. Through prioritization, PRISM extracts high-quality data to learn task-relevant skills effectively. By addressing the impact of noise, our method ensures stable skill learning and achieves superior performance in long-horizon tasks, even with noisy and sub-optimal data.
Machine Learning, ICML
@@ 1 Introduction
Skill-based approaches address these challenges by breaking down long state-action sequences into reusable skills, facilitating hierarchical decision-making and enhancing efficiency in complex tasks (Pertsch et al., 2021, 2022; Shi et al., 2023). Among these, SPiRL (Pertsch et al., 2021) defines skills as temporal abstractions of actions, employing them as low-level policies within a hierarchical framework to achieve success in long-horizon tasks. SiMPL (Nam et al., 2022) builds on this by extending skill learning to meta-RL, using offline expert data to train skills and a context-based high-level policy for task-specific skill selection. Despite these advancements, such methods are highly susceptible to noisy offline demonstrations, which can destabilize skill learning and reduce reliability. In real-world settings, noise often stems from factors like infrastructure aging or environmental perturbations, underscoring the importance of addressing this challenge (Brys et al., 2015; Chae et al., 2022; Yu et al., 2024).
While other RL domains have explored methods to handle noisy demonstrations (Sasaki & Yamashina, 2020; Mandlekar et al., 2022), skill-based learning has largely remained underexplored in this context. To tackle this gap, we propose Prioritized Refinement for Skill-Based Meta-RL (PRISM), a robust framework that incorporates two key contributions: (1) A prioritized skill refinement framework that leverages an exploration policy to discover useful trajectories near noisy offline data, extracting high-quality data through prioritization and employing disentangled skill learning to integrate skills from both online and offline datasets. (2) Maximum return relabeling, a novel technique for evaluating noisy offline trajectories by relabeling their returns based on task relevance, ensuring the selection of high-quality data. By dynamically balancing and prioritizing data from online and offline datasets, PRISM ensures stable and effective skill learning even in noisy environments. These innovations significantly enhance the robustness and generalizability of skill-based meta-RL, delivering reliable performance in real-world noisy scenarios.
Fig. 1 illustrates how the proposed algorithm learns effective skills from noisy demonstrations in the Maze2D environment, where the agent starts at a designated point and must reach an endpoint for each task. Fig. 1(a) shows noisy offline trajectories, which fail to produce effective skills when used directly. In contrast, Fig. 1(b) demonstrates how the prioritized refinement framework uses the exploration policy to navigate near noisy trajectories, identifying paths critical for solving long-horizon tasks and refining useful skills through prioritization. Finally, Fig. 1(c) shows how the high-level policy applies these refined skills to successfully solve unseen tasks. These results highlight the method’s ability to refine and prioritize skills from noisy datasets, ensuring stable learning and enabling the resolution of long-horizon tasks in unseen environments. This paper is organized as follows: Section 3 provides an overview of meta-RL and skill learning, Section 4 details the proposed framework, and Section 5 presents experimental results showcasing the framework’s robustness and effectiveness, along with an ablation study of key components.
@@ 2 Related Works
Skill-based Reinforcement Learning: Skill-based RL has gained traction for tackling complex tasks by leveraging temporally extended actions. Researchers have proposed information-theoretic approaches to discover diverse and predictable skills (Gregor et al., 2016; Eysenbach et al., 2018; Achiam et al., 2018; Sharma et al., 2019), with recent work improving skill quality through additional constraints and objectives (Strouse et al., 2022; Park et al., 2022, 2023; Hu et al., 2024). In offline scenarios, approaches focus on learning transferable behavior priors and hierarchical skills from demonstration data (Pertsch et al., 2021, 2022; Xu et al., 2022; Shi et al., 2023). Building upon these foundations, various skill-based meta-RL approaches have been developed, from hierarchical and embedding-based methods (Nam et al., 2022; Chien & Lai, 2023; Cho & Sun, 2024) to task decomposition strategies (Yoo et al., 2022; He et al., 2024) and unsupervised learning frameworks (Gupta et al., 2018; Jabri et al., 2019; Shin et al., 2024).
Hierarchical Frameworks: Hierarchical approaches in RL have been pivotal for solving long-horizon tasks, where various methods have been proposed including goal-conditioned learning (Levy et al., 2019; Li et al., 2019; Gehring et al., 2021), and option-based frameworks (Bacon et al., 2017; Riemer et al., 2018; Barreto et al., 2019; Araki et al., 2021). Recent advances in goal-conditioned RL have focused on improving sample efficiency (Robert et al., 2024), offline learning (Park et al., 2024), and robust state representations (Yin et al., 2024). The integration of hierarchical frameworks with meta-RL has shown significant potential for rapid task adaptation and complexity handling (Frans et al., 2018; Fu et al., 2020a, 2023). Recent work has demonstrated that hierarchical architectures in meta-RL can provide theoretical guarantees for learning optimal policies (Chua et al., 2023) and achieve efficient learning through transformer-based architectures (Shala et al., 2024).
@@ 3 Background
3.1 Meta-Reinforcement Learning Setup
In meta-RL, each task $\mathcal{T}$ follows a task distribution $p(\mathcal{T})$ and is defined as a Markov Decision Process (MDP) environment $\mathcal{M}^{\mathcal{T}}=\bigl{(}\mathcal{S},\mathcal{A},R^{\mathcal{T}},P^{\mathcal{T}},\gamma\bigr{)}$, where $\mathcal{S}\times\mathcal{A}$ represents the shared state-action space, $R^{\mathcal{T}}$ denotes the reward function, $P^{\mathcal{T}}$ is the state transition probability, and $\gamma$ is the discount factor. At each time $t$, the agent selects an action $a_{t}$ using the policy $\pi$, receives a reward $r_{t}=R^{\mathcal{T}}(s_{t},a_{t})$, and transitions to the next state $s_{t+1}~{}\sim P^{\mathcal{T}}(\cdot|s_{t},a_{t})$. The objective of meta-RL is to train a policy $\pi$ to maximize the return $G=\sum_{t}\gamma^{t}r_{t}$ for the training task set $\mathcal{M}_{\mathrm{train}}$ while enabling the policy to rapidly adapt to unseen test task set $\mathcal{M}_{\mathrm{test}}$, where $\mathcal{M}_{\mathrm{train}}\cap\mathcal{M}_{\mathrm{test}}=\emptyset$.
3.2 Offline Dataset and Skill Learning
To address long-horizon tasks, skill learning from an offline dataset $\mathcal{B}_{\text{off}}:=\{\tilde{\tau}_{0:H}\}$ is considered, which comprises sample trajectories $\tilde{\tau}_{t:t+k}:=(s_{t},a_{t},\cdots,s_{t+k})$ without reward information, where $H$ is the episode length. The dataset $\mathcal{B}_{\text{off}}$ is typically collected through human interactions or pretrained policies. Among various skill learning methods, SPiRL (Pertsch et al., 2021) focuses on learning a reusable low-level policy $\pi_{l}$, using $q(\cdot|\tilde{\tau}_{t:t+H_{s}})$ as a skill encoder to extract the skill latent $z$ by minimizing the following loss function:
$\vspace{-.5em}\mathbb{E}_{\begin{subarray}{c}\tilde{\tau}_{t:t+H_{s}}\sim\mathcal{B}_{\text{off}},\\ z\sim q(\cdot|\tilde{\tau}_{t:t+H_{s}})\end{subarray}}\left[\mathcal{L}(\pi_{l},q,p,z)\right],$ (1)
[…]
3.3 Skill-based Meta-Reinforcement Learning
SiMPL (Nam et al., 2022) integrates skill learning into meta-RL by utilizing an offline dataset of expert demonstrations across various tasks. The skill policy $\pi_{l}$ is trained via SPiRL, while a task encoder $q_{e}$ extracts the task latent $e^{\mathcal{T}}\sim q_{e}$ using the PEARL (Rakelly et al., 2019) framework, a widely-used meta-RL method. During meta-training, the high-level policy $\pi_{h}(z|s,e^{\mathcal{T}})$ selects a skill latent $z$ and executes the skill policy $\pi_{l}(a|s,z)$ for $H_{s}$ time steps, optimizing $\pi_{h}$ to maximize the return for each task $\mathcal{T}$ as:
@@ 4 Methodology
4.1 Motivation: Challenges in Direct Skill Learning with Noisy Demonstrations

### v3

# Self-Improving Skill Learning for Robust Skill-based Meta-Reinforcement Learning

@@ Abstract
Meta-reinforcement learning (Meta-RL) facilitates rapid adaptation to unseen tasks but faces challenges in long-horizon environments. Skill-based approaches tackle this by decomposing state-action sequences into reusable skills and employing hierarchical decision-making. However, these methods are highly susceptible to noisy offline demonstrations, leading to unstable skill learning and degraded performance. To address this, we propose Self-Improving Skill Learning (SISL), which performs self-guided skill refinement using decoupled high-level and skill improvement policies, while applying skill prioritization via maximum return relabeling to focus updates on task-relevant trajectories, resulting in robust and stable adaptation even under noisy and suboptimal data. By mitigating the effect of noise, SISL achieves reliable skill learning and consistently outperforms other skill-based meta-RL methods on diverse long-horizon tasks.
@@ 1 Introduction
Skill-based approaches address these challenges by breaking down long state-action sequences into reusable skills, facilitating hierarchical decision-making and enhancing efficiency in complex tasks (Pertsch et al., 2021; 2022; Shi et al., 2023). Among these, SPiRL (Pertsch et al., 2021) defines skills as temporal abstractions of actions, employing them as low-level policies within a hierarchical framework to achieve success in long-horizon tasks. SiMPL (Nam et al., 2022) builds on this by extending skill learning to meta-RL, using offline expert data to train skills and a context-based high-level policy for task-specific skill selection. Despite these advancements, such methods are highly susceptible to noisy offline demonstrations, which can destabilize skill learning and reduce reliability. In real-world settings, noise often stems from factors like infrastructure aging or environmental perturbations, making it crucial to design methods that remain robust under such conditions (Brys et al., 2015; Chae et al., 2022; Yu et al., 2024a).
While noisy demonstration handling has been explored in other RL settings (Sasaki & Yamashina, 2020; Mandlekar et al., 2022), skill-based meta-RL has largely overlooked this challenge. We identify a critical failure mode: when offline data are suboptimal, the skill library becomes corrupted, and this degradation propagates to the high-level policy, ultimately harming adaptation performance. To address this, we propose Self-Improving Skill Learning (SISL), a robust skill-based meta-RL framework with two key contributions: (1) Decoupled skill self-improvement, achieved through a dedicated improvement policy that perturbs trajectories near the offline data distribution, discovers higher-quality rollouts, and supervises its own updates via a prioritized online buffer. This process progressively denoises the skill library while preserving stability. (2) Skill prioritization via maximum return relabeling, which evaluates offline trajectories with a learned reward model, assigns task-relevant hypothetical returns, and reweights them through a softmax prioritization scheme. This suppresses noisy or irrelevant samples and focuses skill updates on the most beneficial trajectories for downstream adaptation. Together, these components dynamically balance offline and online data contributions, yielding a progressively cleaner skill library and accelerating meta-RL convergence. To our knowledge, SISL is the first framework to explicitly address suboptimal demonstrations in skill-based meta-RL through both exploration-guided refinement and principled relabeling, significantly improving robustness and generalization of skill-learning in real-world noisy scenarios.
Fig. 1 illustrates how the proposed algorithm learns effective skills from noisy demonstrations in the Maze2D environment, where the agent starts at a designated point and must reach an endpoint for each task. Fig. 1(a) shows noisy offline trajectories, which fail to produce effective skills when used directly. In contrast, Fig. 1(b) demonstrates how the prioritized refinement framework uses the improvement policy to navigate near noisy trajectories, identifying paths critical for solving long-horizon tasks and refining useful skills through prioritization. Finally, Fig. 1(c) shows how the high-level policy applies these refined skills to successfully solve unseen tasks. These results highlight the method’s ability to refine and prioritize skills from noisy datasets, ensuring stable learning and enabling the resolution of long-horizon tasks in unseen environments. This paper is organized as follows: Section 3 provides an overview of meta-RL and skill learning, Section 4 details the proposed framework, and Section 5 presents experimental results showcasing the framework’s robustness and effectiveness, along with an ablation study of key components.
@@ 2 Related Works
Skill-based Reinforcement Learning: Skill-based RL has gained traction for tackling complex tasks by leveraging temporally extended actions. Researchers have proposed information-theoretic approaches to discover diverse and predictable skills (Gregor et al., 2016; Eysenbach et al., 2018; Achiam et al., 2018; Sharma et al., 2019), with recent work improving skill quality through additional constraints and objectives (Strouse et al., 2022; Park et al., 2022; 2023; Hu et al., 2024). In offline scenarios, approaches focus on learning transferable behavior priors and hierarchical skills from demonstration data (Pertsch et al., 2021; 2022; Shi et al., 2023; Xu et al., 2022; Kipf et al., 2019; Rana et al., 2023). Building upon these foundations, various skill-based meta-RL approaches have been developed, from hierarchical and embedding-based methods (Nam et al., 2022; Chien & Lai, 2023; Cho & Sun, 2024) to task decomposition strategies (Yoo et al., 2022; He et al., 2024) and unsupervised frameworks (Gupta et al., 2018; Jabri et al., 2019; Shin et al., 2024). For clarity, we use skill in the canonical skill-based RL sense: a latent over fixed-length action sequences. Therefore, studies (Yu et al., 2024b; Wu et al., ) that lack explicit skill learning are not categorized here.
@@
Hierarchical Frameworks: Hierarchical approaches in RL have been pivotal for solving long-horizon tasks, where various methods have been proposed including goal-conditioned learning (Levy et al., 2019; Li et al., 2019; Gehring et al., 2021), and option-based frameworks (Bacon et al., 2017; Riemer et al., 2018; Barreto et al., 2019; Araki et al., 2021). The integration of hierarchical frameworks with meta-RL has shown significant potential for rapid task adaptation and complexity handling (Frans et al., 2018; Fu et al., 2020a; 2023). Recent work has demonstrated that hierarchical architectures in meta-RL can provide theoretical guarantees for learning optimal policies (Chua et al., 2023) and achieve efficient learning through transformer-based architectures (Shala et al., 2024). Recent advances in goal-conditioned RL have focused on improving sample efficiency (Robert et al., 2024), state representation (Yin et al., 2024), and offline-to-online RL (Park et al., 2024; Schmidt et al., 2025). Offline-to-online RL assumes reward-annotated offline datasets to pretrain the policy based on RL before online fine-tuning. In contrast, our setting provides only reward-free offline data for skill learning, making direct application of offline-to-online RL infeasible and clearly distinguishing our approach.
@@ 3 Background
Meta-Reinforcement Learning Setup: In meta-RL, each task $\mathcal{T}$ is sampled from a distribution $p(\mathcal{T})$ and defined as an MDP environment $\mathcal{M}^{\mathcal{T}}=\bigl(\mathcal{S},\mathcal{A},R^{\mathcal{T}},P^{\mathcal{T}},\gamma\bigr)$, where $\mathcal{S}\times\mathcal{A}$ represents the state-action space, $R^{\mathcal{T}}$ is the reward function, $P^{\mathcal{T}}$ denotes the state transition probability, and $\gamma$ is the discount factor. At each step $t$, the agent selects an action $a_{t}$ via the policy $\pi$, receives a reward $r_{t}:=R^{\mathcal{T}}(s_{t},a_{t})$, and transitions to $s_{t+1}~\sim P^{\mathcal{T}}(\cdot|s_{t},a_{t})$. The goal of meta-RL is to train $\pi$ to maximize the return $G=\sum_{t}\gamma^{t}r_{t}$ on the training task set $\mathcal{M}_{\mathrm{train}}$ while enabling rapid adaptation to unseen test tasks in $\mathcal{M}_{\mathrm{test}}$, where $\mathcal{M}_{\mathrm{train}}\cap\mathcal{M}_{\mathrm{test}}=\emptyset$.
Offline Dataset and Skill Learning: To address long-horizon tasks, skill learning from an offline dataset $\mathcal{B}_{\text{off}}:=\{\tilde{\tau}_{0:H}\}$ is considered, which comprises sample trajectories $\tilde{\tau}_{t:t+k}:=(s_{t},a_{t},\cdots,s_{t+k})$ without reward information, where $H$ is the episode length. The dataset $\mathcal{B}_{\text{off}}$ is typically collected through human interactions or pretrained policies. Among various skill learning methods, SPiRL (Pertsch et al., 2021) focuses on learning a reusable low-level policy $\pi_{l}$, using $q(\cdot|\tilde{\tau}_{t:t+H_{s}})$ as a skill encoder to extract the skill latent $z$ by minimizing the following loss function:
$\mathbb{E}_{\begin{subarray}{c}\tilde{\tau}_{t:t+H_{s}}\sim\mathcal{B}_{\text{off}},\\ z\sim q(\cdot|\tilde{\tau}_{t:t+H_{s}})\end{subarray}}\left[\mathcal{L}(\pi_{l},q,p,z)\right],$ (1)
where $\mathcal{L}(\pi_{l},q,p,z):=-\sum^{t+H_{s}-1}_{k=t}\log\pi_{l}(a_{k}|s_{k},z)+\lambda^{\text{kld}}_{l}\mathcal{D}_{\text{KL}}\left(q||\mathcal{N}(\mathbf{0},\mathbf{I})\right)+\mathcal{D}_{\text{KL}}(\lfloor q\rfloor||p)$, $H_{s}$ is the skill length, $\lambda^{\text{kld}}_{l}$ is the coefficient for KL divergence (KLD) $\mathcal{D}_{\text{KL}}$, $\lfloor\cdot\rfloor$ is the stop gradient operator, and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represents a Normal distribution with mean $\bm{\mu}$ and covariance matrix $\mathbf{I}$. Here, $p(z|s_{t})$ is the skill prior to obtain the skill distribution $z$ for a given state $s_{t}$ directly. Using the learned skill policy $\pi_{l}$, the high-level policy $\pi_{h}$ is trained within a hierarchical framework using RL methods.
Skill-based Meta-Reinforcement Learning: SiMPL (Nam et al., 2022) integrates skill learning into meta-RL by utilizing an offline dataset of expert demonstrations across various tasks. The skill policy $\pi_{l}$ is trained via SPiRL, while a task encoder $q_{e}$ extracts the task latent $e^{\mathcal{T}}\sim q_{e}$ using the PEARL (Rakelly et al., 2019) framework, a widely-used meta-RL method. During meta-training, the high-level policy $\pi_{h}(z|s,e^{\mathcal{T}})$ selects a skill latent $z$ and executes the skill policy $\pi_{l}(a|s,z)$ for $H_{s}$ time steps, optimizing $\pi_{h}$ to maximize the return for each task $\mathcal{T}$ as:
@@ 4 Methodology
4.1 Motivation: Toward Robust Skill Learning under Noisy Demonstrations